OpenStudy (anonymous):

the center of a circle is at (6,-7) and the diameter of a circle is 22. Which of the following is the equation of the cirlce/

OpenStudy (anonymous):

Please write equations of circle here

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2.~~~(h,k)~~is~~center,~~~~r=radius\]

OpenStudy (anonymous):

a. (x-6)2 + (y+7)2=11 b.(x-6)2 + (y+7)2=11 c. (x+6)2 + (y-7)2=121 d. (x-6)2 +9y+7)2=121

OpenStudy (anonymous):

okay now we need to use above standard equation of circle Can you try to find out radius from diameter ?

OpenStudy (anonymous):

can you plug in your points into the formula I have posted ? \(\large\color{red}{ \bf tip: }\) x--a=x+a

OpenStudy (anonymous):

Use radius = diameter / 2 and we are having diameter = 22 so radius = --- ?

OpenStudy (anonymous):

im getting confused .. okay 6 is x and -7 is y? right

OpenStudy (anonymous):

Here we need to plug ( h , k ) ( 6 , -7 ) into equation of circle can you try ?